If 93 is added to a number, the result is 41 less than three times the number

Found 2 solutions by Fombitz, Horlaye:
Answer by Fombitz(32388)
Answer by Horlaye(27)

Let the unknown number be 'm'
therefore, we have
93+m = 3m-41
collect like terms
3m-m = 93+41
2m = 134
m = 134/2
m = 67
Hence, the unknown number is 67